OpenStudy (anonymous):

find a polynomial function of minimum degree whose zeros are 0,1,-2

OpenStudy (tkhunny):

There are three, so minimum degree will be 3. Each zero results in one factor. A zero of 1 creates a factor of (x-1) Go!

OpenStudy (anonymous):

how can i solve it by using factor theorem

OpenStudy (anonymous):

the factors are \[(x-r_1)(x-r_2)(x-r_2)\]

OpenStudy (anonymous):

where \(r_i\) are the roots (zeros, whatever)

OpenStudy (anonymous):

write your polynomial in factored form, then multiply it out as @tkhunny said, one of the factors is \((x-1)\)

OpenStudy (anonymous):

what is the factor form of 0

OpenStudy (anonymous):

\(x-0\) which on planet earth is written always as just \(x\)

OpenStudy (anonymous):

this one is always confusing most people get the \((x-1)\) and the \((x+2)\) but get confused by the \(0\) part

OpenStudy (anonymous):

then multiply \[x(x-1)(x+2)\] and you are done

OpenStudy (anonymous):

yeah I already get the 1 and -2 but i don't understand the 0

OpenStudy (anonymous):

so 0 wont be included anymore, @satellite73 ?

OpenStudy (kc_kennylau):

(x-0) is equivalent to x @yuihime

OpenStudy (anonymous):

@kc_kennylau , as far as i know, there isn't -0

OpenStudy (kc_kennylau):

(x-0)(x-1)(x+2) = x(x-1)(x+2) @yuihime

OpenStudy (anonymous):

@kc_kennylau thank you so much

OpenStudy (tkhunny):

We still don't have a function: f(x) = ???

OpenStudy (anonymous):

@satellite73 thanks for your answer

OpenStudy (kc_kennylau):

Well I think that is enough for the answer @tkhunny

OpenStudy (tkhunny):

No, it's not. Look at the problem statement. "find a polynomial function " You have not satisfied the requirements if you write only "x(x-1)(x+2)"

OpenStudy (anonymous):

someone in a bad mood tonight? how about \[f(x)=x(x-1)(x+2)\]?

OpenStudy (xapproachesinfinity):

true, you have to FOIl to find the polynomial

OpenStudy (anonymous):

how to you foil three terms?

OpenStudy (tkhunny):

"FOIL" doesn't mean anything. It was fine as soon as we got the "f(x) ="

OpenStudy (tkhunny):

Yes! Down with Reynolds Wrap.

OpenStudy (anonymous):

so by using remainder theorem, how do you solve \[x(x-1)(x+2)\]

OpenStudy (tkhunny):

You don't "solve" an expression. You will need an equation.

OpenStudy (anonymous):

@tkhunny so how do you do it?

OpenStudy (tkhunny):

I'm not sure what the question is. Please state the "Remainder Theorem".

OpenStudy (anonymous):

Remainder Theorem if a polynomial P(x) is divided by x-c, where c is a real number, then the remainder is P(c) @tkhunny

OpenStudy (tkhunny):

Okay, so now what's the question?

OpenStudy (anonymous):

divide \[x ^{3}+x ^{2}+x+1\] by x+1

OpenStudy (tkhunny):

How's your synthetic division?

OpenStudy (anonymous):

well, it can be solve by synthetic as well,

OpenStudy (tkhunny):

Not "as well". That IS the solution. If \(p(x) = x^{3}+x^{2}+x+1\) And we wish to divide by x+1, we get.  Thus, \(\dfrac{p(x)}{x+1} = x^{2} + 1\) with no remainder.